To find the present value of a cash flow expected to be paid or received in the future, you will the future value cash flow bNow that youve thought about the decision rule that should be applied to your decision, apply it to the following security oDROP DOWN OPTIONS AT TOP ARE: MULTIPLY OR DIVIDE
Answer #1

Present calue of future cash flow or annuity

PV = CF1/(1+r) + CF2/(1+r)^2 + ....... CFn/(1+r)^n

or in case of future value  

PV = FV / (1+r)^n

PV = present value

FV = future value

CF = cash flow or value

r = rate of interest or return

n = number of years

The answer for the drop down box or the 1st answer will be divide as shown in the above formula

Calculating the value today of $ 12,000 cash flow expected to receive 8 years from now at 6% per annum.

PV = 12000/(1.06)^8

= 7528.94 or 7529 (approx)

The 1st option is the best suited option as we should always consider investing in a securiy if the present value of discounted cash flows is greater or equal to the cost of security .

JING ASSOCIATES

next most attractive investment means apportunity cost of this investment is a return of 3.75 % on any other security .

PV = 6750 / (1.0375)^5

= 5615.17 or 5615 (approx)

you are paying $ 5896 for a security which worths $ 5615 today . The security is over prised .

You should not invest in jing associates .

why ?

because the cost of security is greater than the discounted value of the securitys  future cash flow . (option 1)